OpenStudy (anonymous):

What is the area of the regular heptagon given below? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

Do you have a picture?

OpenStudy (anonymous):

What is the area of the polygon given below? 

OpenStudy (anonymous):

Average the two heights, then multiply by the width

OpenStudy (anonymous):

5*2=10 7*2=14 6*2=12 11*2=22 6*2=12 5*2=10 and then add them all together?

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

80 sq. units